OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

Parth (parthkohli):

Distance formula says: \(\Large \color{MidnightBlue}{\Rightarrow Distance = \sqrt{(x_2 - x_1)^2 + (y_2 -y_1)^2} }\)

Parth (parthkohli):

Wait

OpenStudy (anonymous):

ok aha

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{-8^2 + 6^2} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{64 + 36} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{100} }\) \(\Large \color{MidnightBlue}{\Rightarrow 10 }\) So, 10 units.

OpenStudy (anonymous):

ok i got 8 so either way i was wrong but i did my math wrong so yeahh but thanks!! =)